OpenStudy (anonymous):

Express the integrand as a sum of partial fractions and evaluate the integral. What is the integral from 0 to 1 of x^3/ (x^2+8x+16) dx

OpenStudy (anonymous):

First noting that \(x^2+8x+16=(x+4)^2\), you can use synthetic division to do some preliminary simplification. Below is the work for dividing \(x^3\) by \(x+4\): \[\begin{array}{c|cccc} -4&1&0&0&0\\ &&-4&16&-64\\ \hline &1&-4&16&-64 \end{array}\]which translates to \[\frac{x^3}{x+4}=x^2-4x+16-\frac{64}{x+4}\]Dividing the quotient by \(x+4\) (the non-remainder part is considered below), you have \[\begin{array}{c|cccc} -4&1&-4&16\\ &&-4&32\\ \hline &1&-8&48 \end{array}\]which means \[\frac{x^3}{(x+4)^2}=x-8+\frac{48}{x+4}-\frac{64}{(x+4)^2}\]